The random sample shown below was selected from a normal distribution.

3,6,8,3,8,8

Complete parts a and b.

a. Construct a 99% confidence interval for the population mean μ.

(1.971,10.03)

​(Round to two decimal places as​ needed.)

b. Assume that sample mean x overbar x and sample standard deviation s remain exactly the same as those you just calculated but that are based on a sample of n=25 observations. Repeat part a. What is the effect of increasing the sample size on the width of the confidence​ intervals?The confidence interval is (______,_______)​(Round to two decimal places as​ needed.)
Answer #1

Solution:

Given that

x x2
3 9
6 36
8 64
3 9
8 64
8 64
\sumx = 36 \sumx2 = 246

The sample mean is \bar x

\bar x = \sum x / n

= (3 + 6 + 8+ 3 + 8 + 8 / n )

= 36 /6

= 6

The sample mean is = 6

  The sample standard is S

S = \sqrt{} ( \sum x2 ) - (( \sum x )2 / n ) / 1 -n )

= \sqrt{} ( ( 246 -( 36 )2 / 6 ) 5

=\sqrt{} (246 -216 / 5 )

= \sqrt{} 30 / 5

= \sqrt{} 6

= 2.4

The sample standard is 2.4

a ) n = 6

\bar x = 6

s = 2.4

n = 6

Degrees of freedom = df = n - 1 = 6 - 1 = 5

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2,df = t0.005,5 = 4.031

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 4.031 * (2.4 / \sqrt 6)

= 395

Margin of error = 3.95

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

6 - 3.95 < \mu < 6 + 3.95

2.05 < \mu < 9.95

(2.05,9.95 )

b ) n = 25

\bar x = 6

s = 2.4

n = 25

Degrees of freedom = df = n - 1 = 25- 1 = 24

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2,df = t0.005,24 = 2.797

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.797 * (2.4 / \sqrt 25)

= 1.34

Margin of error = 1.34

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

6 - 1.34 < \mu < 6 + 1.34

4..66 < \mu < 7.34

(4.66, 7.34 )

The effect of increasing the sample size on the width of the confidence​ intervals is